I am confused about this problem I'm not sure if its right.
8/2m+1 - 1/m-2 = 5/2m+1
8(m-2)-1(2m+1)=5(m-2)
8m+16-2m+1=5m-10
6m-17=5m-10
m-7=-10
m=3
I don't think I'm getting the lcd right or somethin

Answer by sarah_adam(201)

8/2m+1 - 1/m-2 = 5/2m+1
8(m-2)-1(2m+1)=5(m-2)
8m - 16 - 2m - 1 = 5m -10
6m - 17 = 5m -10
6m - 5m = 17-10
m = 7